Last year, X Company sold 66,000 units of its only product for $18.00 each. Total costs were as follows:

Cost of goods sold
   Variable $415,800
   Fixed 121,440
Selling and administrative
   Variable $94,380
   Fixed 78,540


At the end of the year, a company offered to buy 4,730 units of the product but only for $11.00 each. X Company had the capacity to produce the additional units, and even though there would have been no additional selling and administrative costs, it rejected the offer.

5. If X Company had accepted the special order, firm profits would have increased by?


7. X Company's marketing manager felt that in order for regular sales to continue at 66,000 units this year, the company would have had to reduce the regular selling price this year to $17.62. The result of this price reduction would have been to reduce X Company's profits this year by?6. X Company's production manager thought that the special order units might have required additional direct material and direct labor costs per unit of $0.86 and $0.27, respectively, and the rental of special equipment for $2,000. The combined effect of these changes would have been to reduce the special order profit by how much?
Answer #1
5
Variable cost of goods sold 6.30 =415800/66000
Incremental revenue 52030 =4730*11
Incremental costs:
Variable cost of goods sold 29799 =4730*6.30
Total Incremental costs 29799
Profits increase 22231
6
Increase in cost of goods sold 5345 =4730*(0.86+0.27)
Rental of Equipment 2000
Total decrease in special order profit 7345
7
Reduction in Company's Profits 25080 =66000*(18-17.62)